OpenStudy (anonymous):

Okay this is super simple, but for some reason i keep getting the wrong answer. Help? If f(x) = 2x + 3 and g(x) = 4x - 5, what is f(x) ◦ g(x)?

OpenStudy (anonymous):

I keep getting 8x+2x-15

OpenStudy (anonymous):

Sorry i meant 8x^2

OpenStudy (anonymous):

Ok I see what you did! You are supposed plug g(x) into f(x) but remember you need to put parentheses when you put it in! f(x)=2(4x-5)+3 f(x)=8x-10+3 f(x)=8x-7, get it?

OpenStudy (anonymous):

Oooooooooh that's right, I forgot that and just foiled, haha ok thanks!

OpenStudy (anonymous):

If the dot were filled in it would mean multiplication then you would be right, maybe that's where you got mixed up :P